OpenStudy (anonymous):

4y+15=6y-11 show work

OpenStudy (anonymous):

\[4y+15=6y-11\]subtract the 4y from both sides\[26=2y\]add 11 to both sides\[13=y\]\[y=13\]

OpenStudy (anonymous):

4y-6y=-11-15 -2y=-26 y=13

OpenStudy (anonymous):

the ading eleven happens with the subtracting of 4y lol sorry my bad

OpenStudy (anonymous):

4y+15=6y-11 -15 -15 ________________ 4y=6y-26 -6y -6y __________ -2y=-26 y=13

OpenStudy (anonymous):

then you divide by 2 :)

OpenStudy (anonymous):